A water pump lifts water from a level 10 m below the ground.

Water is pumped at a rate of 30kg/minute with negligible velocity and friction. Calculate the minimum horsepower the engine should have for this



1
Expert's answer
2022-03-23T13:53:00-0400

Given :-Mass, pump Flow rate = 30 kg / min. = 30 kg / 60 s. = 0.5 kg/s.

Height h = 10 m.

we know, pump water has negligible velocity , then it's K.E is zero.

Pump work to increase the level of water above 10 m.

weight of the water = mg

= 0.5 × 9.8 = 4.9 N.

power = mgh

P= 4.9 × 10 m

P= 49 W.

1 hp = 746 W.

minimum power = 49 / 746

 =0.066= 6.6 × 10⁻² hp.

therefore, the minimum horsepower is 6.6 x 10⁻² hp.